# Explaining Similarity in Vision–Language Encoders with Weighted Banzhaf Interactions

### v1

@@ 5.1 FIxLIP outperforms baselines as measured with insertion/deletion curves
Figure 3: Insertion/deletion curves for CLIP (ViT-B/32) on MS COCO. AID score (higher is better) for FIxLIP against alternative explanation methods, where a random baseline scores 0. The y-axis is normalized between the model’s prediction on the original input ($100\%$) and the fully removed one ($0\%$), where negative values denote that the model is predicting the image–text inputs are unsimilar. It means the similarity prediction on a partially masked input is smaller than the prediction on the fully masked input. Methods like Grad-ECLIP and XCLIP fail to recover nonlinear rankings of important tokens, while our method faithfully recovers the optimal subset explanation. Extended results for CLIP (ViT-B/16) and SigLIP-2 (ViT-B/32) are in Figures 10 & 11.
@@ 5.2 First-order attribution methods fail to pass a sanity check with a pointing game
Table 1 shows PGR results for different explanation methods, where each metric value is an average ($\pm$ sd.) over 500 inputs. We observe that first-order methods like Banzhaf values fail to discriminate between multiple objects in an image. Saliency maps can only highlight the important part of image–text inputs without disentangling the complex relationship between each separate word in a caption and the corresponding image regions (Appendix Figure 8). Second-order methods like FIxLIP and XCLIP pass our proposed sanity check; the latter scores a bit higher as it specializes in approximating cross-modal interactions. The general conclusion is consistent for CLIP (ViT-B/16). We further analyze the potential application of a pointing game to evaluate different models in Section 5.6.
Table 1: Pointing game recognition for CLIP (ViT-B/32) on ImageNet-1k. PGR score (higher is better) for FIxLIP against alternative explanation methods, where a random baseline scores $0.25$. The 1st-order methods like Grad-ECLIP and Shapley values fail to distinguish between multiple objects at once, while 2nd-order methods faithfully recover the appropriate explanation (up to the pointing game’s irreducible non-optimality). Extended results for CLIP (ViT-B/16) are in Table 4.
@@
GAME (Chefer et al., 2021) $61_{\pm.12} .$$43_{\pm.03} .$$33_{\pm.02} .$$28_{\pm.01}$
Grad-ECLIP (Zhao et al., 2024) $68_{\pm.15} .$$45_{\pm.04} .$$33_{\pm.02} .$$28_{\pm.01}$
@@ 5.3 FIxLIP recovers a faithful decomposition of the similarity function
Figure 4 demonstrates $p$-faithfulness results for different explanation methods, where each boxplot represents statistics aggregated from 1000 inputs. We confirm that FIxLIP optimizes faithfulness, which depends on the parameter $p$. Weighted Banzhaf interactions allow for precise controllability of the faithfulness optimization, while Shapley interactions perform well on average. Alternative attribution methods compute explanations unfaithful to the explained similarity function ($\mu_{\mathrm{corr}}\approx 0.5$). Extended results for different models and $p$ values are given in Appendix D.
@@ 5.4 On the computational efficiency of the FIxLIP cross-modal estimator
Figure 5 demonstrates that the FIxLIP cross-modal estimator achieves over $20\times$ speedup when considering model inference time, i.e. game evaluations, and about $5\times$ speedup when accounting for the entire explanation pipeline implemented in Python. We did not necessarily optimize the latter, i.e. subset sampling, weighted least squares optimization, and other processing steps, expecting that the visible gap between the explanation and game time could be further decreased. For context, related first-order attribution methods take about 1 second to compute; see Table 5 in (Zhao et al., 2024), but note that it could be only for an image explanation without text attribution.
@@ 5.5 Visual explanation of vision–language encoders
We have already established that FIxLIP delivers state-of-the-art faithfulness performance across three diverse metrics. One of the explanations’ application is to interpret the model’s output function. Figure 6 demonstrates three types of visualizations that we envision can be used to broaden our understanding of VLEs. In the interest of space, we provide further interesting examples in Appendix E. Specifically, we provide a comprehensive guide on interpreting interaction explanations like FIxLIP in Appendix E.1. Figures 8 & 9 further compare FIxLIP to baselines in the context of evaluation metrics. Figure 16 shows a comparison between FIxLIP of CLIP and SigLIP-2 (ViT-B/32), while Figure 17 shows a comparison between the ViT-B/32 and ViT-B/16 versions of CLIP.
@@ 5.6 Model-agnostic comparison of different vision–language encoder architectures
We use FIxLIP to compare different VLE architectures with PGR in Table 2. Surprisingly, SigLIP-2 can be more faithfully explained with cross-modal interactions than CLIP, for both model sizes. We observe that, in general, smaller models (ViT-B/32) can be more faithfully explained than larger ones (ViT-L/16), which is consistent with prior work (Bordt and von Luxburg, 2023; Enouen and Liu, 2025). For more visual comparisons, see Appendix E.3.
@@ 6 Discussion
Limitations and future work. Our work faces three limitations with clear paths for future developments. Although FIxLIP allows scaling to larger models with hundreds of players and efficiently computing a million model inferences, improvements could be made to its practical implementation. Specifically, we envision exploring the use of sparse linear regression, applying the Archipelago framework (Tsang et al., 2020) to filter interactions for the approximation, and a non-greedy algorithm for a closer-to-optimal subset selection. Second, the visual properties and usability of FIxLIP should be further studied as a potential direction in human-computer interaction research (Rong et al., 2024), see e.g. a comprehensive user study of cross-modal interaction explanations by Liang et al. (2023). Finally, our work is restricted to second-order interaction explanations, while efficiently approximating higher-order interactions (also in tri-modal settings beyond LIP models) becomes an interesting challenge to overcome.
@@
We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within the computational grant no. PLG/2025/018330. Hubert Baniecki was supported from the state budget within the Polish Ministry of Education and Science program “Pearls of Science” project number PN/01/0087/2022. Maximilian Muschalik and Fabian Fumagalli gratefully acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation): TRR 318/1 2021 – 438445824.
Appendix for “Explaining Similarity in Vision–Language Encoders with Weighted Banzhaf Interactions”
@@ A.1 Proof of Theorem˜1
The approximated game $\hat{\nu}_{\mathbf{e}}$ is a 2-additive game, and hence its Möbius transform is restricted up to second-order interactions [Grabisch, 2016]. Moreover, the Möbius transform of the approximated game $\hat{a}^{\textsc{FIxLIP}}$ satisfies
@@
where $i,j\in N_{\mathcal{I}}\cup N_{\mathcal{T}}$ and $i\neq j$. The approximated game $\hat{\nu}$ is by construction a 2-additive game, and thus all higher-order Möbius coefficients are zero [Grabisch, 2016]. It was further shown by Marichal and Mathonet [2011, Proposition 3] that the weighted Banzhaf values of $\hat{\nu}$ can be computed as the optimal first-order approximation of $\hat{\nu}$. The representation of the best $k$-th order approximation in the Möbius transform was given by Marichal and Mathonet [2011, Proposition 6], and reads for $i\in N_{\mathcal{I}}\cup N_{\mathcal{T}}$, $k=1$, and equal probabilities $p$ as
@@ A.2 Proof of Theorem˜2
We now use the law of total variance and $M_{\mathcal{I}}\perp M_{\mathcal{T}}$ to re-write
@@
and use the similar result for $M_{\mathcal{T}}$ to obtain
@@ B.1 Faithful interaction explanations of LIP models (FIxLIP)
Masking. Following related work using input removal in attribution methods [Hase et al., 2021, Erion et al., 2021, Lundstrom et al., 2022, Fumagalli et al., 2023], we apply simple baseline masking strategies, which were also proven successful based on our empirical results. For vision encoders, we use $\mathbf{0}$ baseline after image normalization to propagate no signal forward. In the case of language encoders, we use a native attention masking mechanism, where we encode active tokens with $1$ and the deleted ones with 0. It is important to leave the beginning/end sequence and padding tokens as is. Note that in theory, the maximum possible context length $n_{\mathcal{T}}$ is finite (e.g. 64, 77) and could be treated as constant across different inputs. However, we never explain these special tokens, treating parts of text inputs as the only features of interest. Alternative implementations of masking worth considering are: removing words from the text inputs, or tokens from tokenized inputs, imputing with the [UNK] token, or [MASK], e.g. in the case of the BiomedCLIP architecture that is based on BERT [Zhang et al., 2024]. Acknowledging the potential influence of the masking approach, as well as position encodings, on the value of an empty input, we check that it remains about constant across different text input lengths in Figure 7.
@@
Two-step filtering approach. For the ViT-B/16 and ViT-L/16 model versions with a larger number of input image tokens (patches), we perform cheap computation of the first-order attribution to prioritize interactions for approximation. We consider two strategies: picking a clique subset of the most interesting tokens, or focusing only on cross-modal interaction for explanation, which we specifically apply in the pointing game recognition (Definition 10, Appendix B.3). In the first strategy, given the desirable clique size $k=72$, we take $k_{\mathcal{T}}\coloneq\max\!\big{(}5,\lceil k\cdot n_{\mathcal{T}}/(n_{\mathcal{T}}+n_{\mathcal{I}})\rceil\big{)}$ text tokens with the highest absolute feature attribution scores, and $k_{\mathcal{I}}\coloneq k-k_{\mathcal{T}}$ image tokens accordingly. Similarly to the above considerations regarding the budget split between the two modalities, we found this to be empirically adequate. Future work can propose optimal strategies for both of these nuances.
Greedy subset selection. Even for the ViT-B/32 model version with fewer than a hundred input tokens (nodes in a graph), finding cliques of the highest and smallest sums with brute force is computationally prohibitive. Still, we want to use these subsets denoted as $M_{\mathbf{e},\mathrm{min},k}$ and $M_{\mathbf{e},\mathrm{max},k}$ in evaluation with insertion/deletion curves (Definition 9) and explanation visualization. Thus, we implemented a simple greedy algorithm starting the search from each (or a subset) of the tokens (nodes). Its general goal is to add consecutive tokens to the subset based on minimizing/maximizing the subset’s value. For the models with a larger number of inputs, e.g. $n_{\mathcal{I}}+n_{\mathcal{T}}=196+30$, the greedy strategy remains applicable; although, may require a few minutes to complete when considering all subset sizes $k=1,\dots,n_{\mathcal{I}}+n_{\mathcal{T}}$.
@@ B.2 Area between the insertion/deletion curves (AID)
Figure 9 describes visually the process of computing insertion/deletion curves and how it can differ between the first-order attribution methods and second-order interaction explanations. Consider the following example shown in (bottom, Ours, 73%), where the faces of a smiling man and child are masked with the masked caption saying “a mana child smiling at ain a —”. Here, the model predicts the input to be dissimilar (below-average similarity), which means going into negative values in our normalized case. Contrarily, in (bottom, Baseline), the unmasked restaurant text token keeps the model’s similarity near the average level. Another interesting phenomenon happens for the insertion curve in Figure 9 (top). Faithfully masking the redundant information with our method causes the model’s similarity to increase above the original prediction. Contrarily, in (top, Baseline), the baseline gradient-based method is unable to faithfully recover such redundant tokens.
@@ B.3 Pointing game recognition (PGR)
Figure 8 shows an example of the pointing game evaluation for interaction explanations like FIxLIP, giving additional context as to why first-order attribution methods fail to pass this desirable sanity check. A saliency map can only highlight important parts of image/text inputs, but whenever more objects/concepts appear in the input, they are unable to disentangle the basic relationship between both modalities. PGR metric measures the ratio of absolute values of “correctly” identified cross-modal interaction terms to total interaction terms. For example, in Figure 8 (2nd row, 2nd column), we sum the positive interactions between “banana”/“cat” image tokens and the banana/cat text tokens, as well as the absolute negative interactions between “tractor”/“ball” image tokens and the banana/cat text tokens. Then, we divide this sum by the total absolute sum of all interactions, which gives a normalized PGR score. In Appendix C.2, we further describe the specific combinations of image and object/class labels used to create our exemplary benchmark in this paper, although the overall methodology should be treated in a generic manner.
@@
Figure 9: Exemplary insertion/deletion curves for two explanations $\mathbf{e}$ computed with FIxLIP and another first-order method. We show five subset visualizations per curve per method. In the case of the baseline method, each consecutive subset needs to include the preceding one, which is not the case for FIxLIP capturing the complex cross-modal and intra-modal interactions. For example, a child with the corresponding face patches (top right corner) gets overvalued by the manpizza tokens appearing jointly with the corresponding image part (child disappears). Similarly, in the deletion curve, the subsets of text and image tokens interchange with each other at around 50% (bottom row).

### v2

@@ 5.1 FIxLIP outperforms baselines as measured with insertion/deletion curves
Figure 3: Insertion/deletion curves for CLIP (ViT-B/32) on MS COCO. AID score (higher is better) for FIxLIP against alternative explanation methods, where a random baseline scores 0. The y-axis is normalized between the model’s prediction on the original input ($100\%$) and the fully removed one ($0\%$), where negative values denote that the model is predicting the image–text inputs are unsimilar. It means the similarity prediction on a partially masked input is smaller than the prediction on the fully masked input. Methods such as Grad-ECLIP and exCLIP fail to recover nonlinear rankings of important tokens, while our method faithfully recovers the optimal subset explanation. Extended results for CLIP (ViT-B/16) and SigLIP-2 (ViT-B/32) are in Figures 10 & 11.
@@ 5.2 First-order attribution methods fail to pass a sanity check with a pointing game
Table 1 shows PGR results for different explanation methods, where each metric value is an average ($\pm$ sd.) over 500 inputs. We observe that first-order methods, e.g. Banzhaf values, fail to discriminate between multiple objects in an image. Saliency maps can only highlight the important part of image–text inputs without disentangling the complex relationship between each separate word in a caption and the corresponding image regions (Appendix Figure 8). Second-order methods such as FIxLIP and exCLIP pass our proposed sanity check; the latter scores slightly higher, as it specializes in approximating cross-modal interactions. The conclusion is consistent for CLIP (ViT-B/16). We further analyze the potential application of a pointing game to evaluate different models in Section 5.6.
Table 1: Pointing game recognition for CLIP (ViT-B/32) on ImageNet-1k. PGR score (higher is better) for FIxLIP against alternative explanation methods, where a random baseline scores $0.25$. First-order methods, such as Grad-ECLIP and Shapley values, fail to distinguish between multiple objects at once, while second-order methods faithfully recover the appropriate explanation (up to the pointing game’s irreducible non-optimality). Extended results for CLIP (ViT-B/16) are in Table 4.
@@
GAME [chefer2021generic] $61_{\pm.12} .$$43_{\pm.03} .$$33_{\pm.02} .$$28_{\pm.01}$
Grad-ECLIP [zhao2024gradeclip] $68_{\pm.15} .$$45_{\pm.04} .$$33_{\pm.02} .$$28_{\pm.01}$
@@ 5.3 FIxLIP recovers a faithful decomposition of the similarity function
Figure 4 demonstrates $p$-faithfulness results for different explanation methods, where each boxplot represents statistics aggregated from 1000 inputs. We confirm that FIxLIP optimizes faithfulness, which depends on the parameter $p$. Weighted Banzhaf interactions allow for precise controllability of the faithfulness optimization, whereas Shapley interactions perform well on average. Alternative attribution methods compute explanations unfaithful to the explained similarity function ($\mu_{\mathrm{corr}}\approx 0.5$). Extended results for different models and $p$ values are given in Appendix D.
@@ 5.4 On the computational efficiency of the FIxLIP cross-modal estimator
Figure 5 demonstrates that the FIxLIP cross-modal estimator achieves over $20\times$ speedup when considering model inference time, i.e. game evaluations, and about $5\times$ speedup when accounting for the entire explanation pipeline implemented in Python. We did not necessarily optimize the latter, i.e. subset sampling, weighted least squares optimization, and other processing steps, expecting that the visible gap between the explanation and game time could be further decreased. For context, related first-order attribution methods take about 1 second to compute; see Table 5 in [zhao2024gradeclip], but note that it could be only for an image explanation without text attribution.
@@ 5.5 Visual explanation of vision–language encoders
We have already established that FIxLIP delivers state-of-the-art faithfulness performance across three diverse metrics. One of the explanations’ applications is to interpret the model’s output function. Figure 6 demonstrates three types of visualizations that we envision can be used to broaden our understanding of VLEs. In the interest of space, we provide further interesting examples in Appendix E. Specifically, we provide a comprehensive guide on interpreting interaction explanations like FIxLIP in Appendix E.1. Figures 8 & 9 further compare FIxLIP to baselines in the context of evaluation metrics. Figure 16 shows a comparison between FIxLIP of CLIP and SigLIP-2 (ViT-B/32), while Figure 17 shows a comparison between the ViT-B/32 and ViT-B/16 versions of CLIP.
@@ 5.6 Model-agnostic comparison of different vision–language encoder architectures
We use FIxLIP to compare different VLE architectures with PGR in Table 2. Surprisingly, SigLIP-2 can be more faithfully explained with cross-modal interactions than CLIP, for both model sizes. We observe that, in general, smaller models (ViT-B/32) can be more faithfully explained than larger ones (ViT-L/16), which is consistent with prior work [bordt2023shapley, enouen2025instashap]. For more visual comparisons, see Appendix E.3.
@@ 6 Discussion
Limitations and future work. Our work faces three limitations, each with a clear path for future development. Although FIxLIP allows scaling to larger models with hundreds of players and efficiently computing a million model inferences, improvements could be made to its practical implementation. Specifically, we envision exploring the use of sparse linear regression, applying the Archipelago framework [tsang2020how] to filter interactions for the approximation, and a non-greedy algorithm for a closer-to-optimal subset selection. Second, the visual properties and usability of FIxLIP should be further studied as a potential direction in human-computer interaction research [rong2024towards], see e.g. a user study of cross-modal interaction explanations by liang2023multiviz. Finally, our work is restricted to second-order interaction explanations, while efficiently approximating higher-order interactions (also in tri-modal settings beyond LIP models) becomes an interesting challenge to overcome.
@@
Code. We provide additional details on reproducibility in the Appendix, as well as the code to reproduce all experiments in this paper is available at https://github.com/hbaniecki/fixlip.
@@
We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within the computational grant no. PLG/2025/018330. Hubert Baniecki was supported from the state budget within the Polish Ministry of Education and Science program “Pearls of Science” project number PN/01/0087/2022. Barbara Hammer, Eyke Hüllermeier, Fabian Fumagalli, and Maximilian Muschalik gratefully acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation): TRR 318/1 2021 – 438445824.
Appendix
@@ A.1 Proof of Theorem˜1
The approximated game $\hat{\nu}_{\mathbf{e}}$ is a 2-additive game, and hence its Möbius transform is restricted up to second-order interactions [grabisch2016set]. Moreover, the Möbius transform of the approximated game $\hat{a}^{\textsc{FIxLIP}}$ satisfies
@@
where $i,j\in N_{\mathcal{I}}\cup N_{\mathcal{T}}$ and $i\neq j$. The approximated game $\hat{\nu}$ is by construction a 2-additive game, and thus all higher-order Möbius coefficients are zero [grabisch2016set]. It was further shown by marichal2011weighted that the weighted Banzhaf values of $\hat{\nu}$ can be computed as the optimal first-order approximation of $\hat{\nu}$. The representation of the best $k$-th order approximation in the Möbius transform was given by marichal2011weighted, and reads for $i\in N_{\mathcal{I}}\cup N_{\mathcal{T}}$, $k=1$, and equal probabilities $p$ as
@@ A.2 Proof of Theorem˜2
We now use the law of total variance and $M_{\mathcal{I}}\perp M_{\mathcal{T}}$ to rewrite
@@
and use a similar result for $M_{\mathcal{T}}$ to obtain
@@ B.1 Faithful interaction explanations of LIP models (FIxLIP)
Masking. Following related work using input removal in attribution methods [hase2021outofdistribution, erion2021improving, lundstrom2022rigorous, fumagalli2023shapiq], we apply simple baseline masking strategies, which were also proven successful based on our empirical results. For vision encoders, we use a $\mathbf{0}$ baseline after image normalization to propagate no signal forward. In the case of language encoders, we use a native attention masking mechanism, where we encode active tokens with $1$ and the deleted ones with 0. It is important to leave the beginning/end sequence and padding tokens as is. Note that in theory, the maximum possible context length $n_{\mathcal{T}}$ is finite (e.g. 64, 77) and could be treated as constant across different inputs. However, we never explain these special tokens, treating parts of text inputs as the only features of interest. Alternative implementations of masking worth considering are: removing words from the text inputs, or tokens from tokenized inputs, imputing with the [UNK] token, or [MASK], e.g. in the case of the BiomedCLIP architecture that is based on BERT [zhang2024biomedclip]. Acknowledging the potential influence of the masking approach, as well as position encodings, on the value of an empty input, we check that it remains about constant across different text input lengths in Figure 7.
@@
Two-step filtering approach. For the ViT-B/16 and ViT-L/16 model versions with a larger number of input image tokens (patches), we perform a cheap computation of the first-order attribution to prioritize interactions for approximation. We consider two strategies: picking a clique subset of the most interesting tokens, or focusing only on cross-modal interaction for explanation, which we specifically apply in the pointing game recognition (Definition 10, Appendix B.3). In the first strategy, given the desirable clique size $k=72$, we take $k_{\mathcal{T}}\coloneq\max\!\big(5,\lceil k\cdot n_{\mathcal{T}}/(n_{\mathcal{T}}+n_{\mathcal{I}})\rceil\big)$ text tokens with the highest absolute feature attribution scores, and $k_{\mathcal{I}}\coloneq k-k_{\mathcal{T}}$ image tokens accordingly. Similarly to the above considerations regarding the budget split between the two modalities, we found this to be empirically adequate. Future work can propose optimal strategies for both of these nuances.
Greedy subset selection. Even for the ViT-B/32 model version with fewer than 100 input tokens (nodes in a graph), finding cliques of the highest and smallest sums using brute force is computationally prohibitive. Still, we want to use these subsets denoted as $M_{\mathbf{e},\mathrm{min},k}$ and $M_{\mathbf{e},\mathrm{max},k}$ in evaluation with insertion/deletion curves (Definition 9) and explanation visualization. Thus, we implemented a simple greedy algorithm starting the search from each (or a subset) of the tokens (nodes). Its general goal is to add consecutive tokens to the subset based on minimizing/maximizing the subset’s value. For models with a larger number of inputs, e.g. $n_{\mathcal{I}}+n_{\mathcal{T}}=196+30$, the greedy strategy remains applicable; however, it may require a few minutes to complete when considering all subset sizes $k=1,\dots,n_{\mathcal{I}}+n_{\mathcal{T}}$.
@@ B.2 Area between the insertion/deletion curves (AID)
Figure 9 describes visually the process of computing insertion/deletion curves and how it can differ between the first-order attribution methods and second-order interaction explanations. Consider the following example shown in (bottom, Ours, 73%), where the faces of a smiling man and child are masked with the masked caption saying “a mana child smiling at ain a —”. Here, the model predicts the input to be dissimilar (below-average similarity), which means it enters negative values in our normalized case. Contrarily, in (bottom, Baseline), the unmasked restaurant text token keeps the model’s similarity near the average level. Another interesting phenomenon happens for the insertion curve in Figure 9 (top). Faithfully masking the redundant information with our method causes the model’s similarity to increase above the original prediction. Contrarily, in (top, Baseline), the baseline gradient-based method is unable to faithfully recover such redundant tokens.
@@ B.3 Pointing game recognition (PGR)
Figure 8 illustrates an example of the pointing game evaluation for interaction explanations like FIxLIP, giving additional context as to why first-order attribution methods fail to pass this desirable sanity check. A saliency map can only highlight important parts of image/text inputs, but whenever more objects/concepts appear in the input, they are unable to disentangle the basic relationship between the two modalities. PGR metric measures the ratio of absolute values of “correctly” identified cross-modal interaction terms to total interaction terms. For example, in Figure 8 (2nd row, 2nd column), we sum the positive interactions between “banana”/“cat” image tokens and the banana/cat text tokens, as well as the absolute negative interactions between “tractor”/“ball” image tokens and the banana/cat text tokens. Then, we divide this sum by the total absolute sum of all interactions, which gives a normalized PGR score. In Appendix C.2, we further describe the specific combinations of image and object/class labels used to create our exemplary benchmark in this paper, although the overall methodology should be treated in a generic manner.
@@
Figure 9: Exemplary insertion/deletion curves for two explanations $\mathbf{e}$ computed with FIxLIP and another first-order method. We show five subset visualizations per curve per method. In the case of the baseline method, each consecutive subset needs to include the preceding one, which is not the case for FIxLIP, capturing the complex cross-modal and intra-modal interactions. For example, a child with the corresponding face patches (top right corner) gets overvalued by the manpizza tokens appearing jointly with the corresponding image part (child disappears). Similarly, in the deletion curve, the subsets of text and image tokens interchange with each other at around 50% (bottom row).